OpenStudy (anonymous):

Find the general term an for the geometric sequence. \[a_{1}=-2\] and \[a_{2}=6\] \[a_{n}=\]

OpenStudy (anonymous):

Divide a2 by a1 and you have the ratio. Then, divide a1 by that to get your initial term a0 and the formula will be an=a0*r^n

OpenStudy (anonymous):

1.5

OpenStudy (anonymous):

@Numb3r1

OpenStudy (anonymous):

No, 6/-2 does not equal 1.5.

OpenStudy (anonymous):

ohhh!! I divided it by -3

OpenStudy (anonymous):

instead

OpenStudy (anonymous):

its -3

OpenStudy (anonymous):

@Numb3r1

OpenStudy (anonymous):

Yup. Now divide -2 by that for the initial term, and your formula will be an=a0*(-3)^n

OpenStudy (anonymous):

i did and i got 1.5

OpenStudy (anonymous):

@Numb3r1

OpenStudy (anonymous):

No, divide -2/-3, not -3/-2. 1/1.5 will give you the same answer.

OpenStudy (anonymous):

0.6

OpenStudy (anonymous):

0.66666666666666 is closer, but to be perfectly correct, 2/3 is best.

OpenStudy (anonymous):

oh!! since it can't be factored it becomes 2/3. DUH!!!

OpenStudy (anonymous):

@Numb3r1 so my answer is 2/3?

OpenStudy (anonymous):

an=a0*r^n an=(2/3)*(-3)^n

OpenStudy (anonymous):

i dont get it :(

OpenStudy (anonymous):

\[\frac{ 2 }{ 3}\left( -3 \right)^{n}\]

OpenStudy (anonymous):

@Numb3r1

OpenStudy (anonymous):

@e.mccormick can you help?

OpenStudy (e.mccormick):

Yah, that was it in the end.